how do i solve these equations? Please show step by step. thanks
Find the roots of x to the second power + 11x= -30?
Find the roots of 3x to the second power + 2x-4=0?
Find the roots of x to the second power -8+17?

Answer by checkley71(8403)

X~2+11X+30=0 OR (X+5)(X+6)=0 OR X=-5 & X=-6
3X~2+2X-4=0 OR USING THE QUADRIC EQUATION SOLUTION (-B+/-SQRT[B~2-4AC])/2A THUS THE VALUES OF X ARE .868517 & -1.535183
X~2-8+17=0 OR X~2+9=0 OR AGAIN USING THE QUADRIC FORMULA WE GET x=3i & -3i